OpenStudy (anonymous):

Solve the inequality. Show your work. |r + 3| ≥ 7

OpenStudy (anonymous):

im bad with these

jimthompson5910 (jim_thompson5910):

If \[\large |r+3|\ge 7\] then \[\large r+3\ge 7 \ \text{or} \ r+3\le -7\]

jimthompson5910 (jim_thompson5910):

solve each inequality